If loudness of sound of a barking dog at 4 m is 40dB ,then loudness at 8m from dog is (assume dog to be point of source) approx.?

Grade:

1 Answers

Arun
5 years ago
loudness of sound is inversely proportional to the square of the distance.
Lesser the  distance between the source and the listener, greater is the loudness of the sound heard by the listener and vice -versa.
 
let new loudness = x
 
hence
 
x/40 = 16/64
 
x = 640/64
 
x = 10dB